OpenStudy (anonymous):

[(51 + 3) − 32] ÷ 9 • 2

OpenStudy (anonymous):

Its kinda stupid but i forgot how to do it..

OpenStudy (theviper):

[(51 + 3) − 32] ÷ 9 • 2 [54-32]/9*2 22/9*2 22/18=??

OpenStudy (anonymous):

1.2? i guess

OpenStudy (anonymous):

Thank you :)

Parth (parthkohli):

@TheViper no.

Parth (parthkohli):

You do it from left to right, so you will have 22/9 * 2 = 44/9